OpenStudy (loser66):

any circle can be written in the form \(x^2 + y^2 + cx + dy +e =0\) Write equation of the circle passes through A (0,1) , B ( 1,3), C ( 2,5). Please, help

imqwerty (imqwerty):

hmm a length method will be to find the equation of AB and BC and then their normal the point where the normals intersect is the centre once we get the center we can find the radius and then we r done

OpenStudy (loser66):

Replace: for A: d + e =-1(1) for B: c + 3d +e = -10 (2) for C: 2c + 5d +e = -29 (3) If I multiple (2) by 2 and subtract (3) from it, I have d +e = 9 hence with (1) , we have no solution. But it is obviously wrong. However, I don't see what is wrong with the logic.

OpenStudy (loser66):

When solving it by matrix, I got c = -5, d= -2, e = 1. But I need know what is wrong with the logic above.

OpenStudy (zarkon):

isn't A, B and C colinear

OpenStudy (amistre64):

A (0,1) , B ( 1,3), C ( 2,5) cx+dy+e=-(x^2+y^2) 0c+1d+1e=-1 1c+3d+1e=-10 2c+5d+1e=-29 by elimination ... -2c-6d-2e= 20 2c+5d+1e=-29 -------------- d+e = 9 rref{{0,1,1},{1,3,1},{2,5,1}} = 1,0,-2 0,1,1 0,0,0 we have a free .... radical? forget the term ... but the matrix is not independant d+e = 9 d+e = 1

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

you got a line there with slope two and y intercept one must be a magic circle

OpenStudy (misty1212):

oh, what @Zarkon said

imqwerty (imqwerty):

lol yes they r collinear xD so circle shuld not exist slope of AB= slope of BC cx

OpenStudy (loser66):

:) Shame on me.!! Thanks all .

OpenStudy (amistre64):

if its a circle, then you are looking at it from the wrong angle :)

OpenStudy (loser66):

But I don't know why I put it on my calculator, it gives me the answer c = -5, d = -2, e =1. :(((

OpenStudy (amistre64):

user error?

OpenStudy (loser66):

I put the numbers in 3 x 4 matrix and hit the rref, the result pops out like this. hehehe.... it is crazy like me.

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=rref {{1%2C3%2C1%2C-10}%2C{2%2C5%2C1%2C-29}%2C{0%2C1%2C1%2C-1}}

OpenStudy (amistre64):

firefox likes to post broken links i spose

OpenStudy (loser66):

is it not that